# Charge collection efficiency of thimble ionization chambers exposed to ultra-high dose per pulse

### v1

@@ Abstract
Background: Commercially available ionization chambers (ICs) exposed to ultra-high dose per pulse (DPP) exhibit deviations from a linear dose response due to volume recombination. In the last years, phenomenological and simulation models have been developed to describe the charge collection efficiency (CCE) focused on parallel-plate ICs.
Purpose: To measure and simulate the CCE and polarity effect of thimble ICs exposed to ultra-high DPPs.
@@
Methods: The response of two PinPoint3D T31022 (PP3D) and two PinPoint T31023 (PP) ICs was investigated experimentally at the national metrology institute of Germany (PTB). The ICs were irradiated using the ultra-high-DPP reference electron beam with an energy of 20 MeV and DPPs between 0.1 Gy up to 9.3 Gy. The bias voltage supplied to the ICs was varied between 200 V up to 500 V. Additionally, the time-resolved signal of the ICs was recorded using an oscilloscope. To simulate the response of the chambers, a novel finite element code capable of simulating 1D and 2D geometries was developed. Three different geometries were considered to describe the investigated ICs: a cylindrical 1D geometry, a simplified 2D geometry and a complete 2D geometry including the conductive guard ring of the ICs.
@@
Results: The experimental data shows that thimble ICs exhibit a pronounced polarization effect when irradiated with high DPP. When the sign of the collected charge it is negative, the CCE is larger than when it is positive. The main part of the contribution to the polarity effect arises from the polarity-dependent charge transport and recombination inside the IC. Based on simulation, the CCE of an ideal thimble IC can be described as a volume weighted-average between the CCE of a cylindrical and a spherical IC. Discrepancies between this weighted rule and a simplified 2D model are below 2.5 % and 0.2 % for the PP3D and the PP geometries, respectively, in terms of CCE at 300 V. When the guard ring is considered in the simulation, the CCE and polarity correction factor are reduced, and differences in CCE with the volume weighted model up to 18 % and 1.7 % for the PP3D and PP are found. Experimental and simulated CCE (polarity effect correction factor) agrees within 1.4 % (7.0 %) and 1.6 % (3.2 %) for the PP3D and PP ICs, respectively. The CCE of a parallel-plate IC and a thimble IC can be related using a simple geometrical rule.
Conclusions: Thimble ICs exposed to ultra-high-DPP exhibit a large polarity effect due to the different distribution and recombination of the charge carriers whether the free electrons drift toward the central or outer electrode. Although the two thimble ICs studied have a similar sensitive volume, the PP shows a greater CCE due to its smaller external radius. A numerical model based on the finite element method is able to satisfactory reproduce the actual CCE for these two chambers. For the PP3D, the inclusion of the guard ring in the simulation geometry is mandatory to obtain accurate results. At large DPPs, thimble ICs should be used with caution due to their large polarity effect.
Contents
I. Introduction
@@
Commercially available ionization chambers (IC) suffer from volume recombination when exposed to ultra-high dose per pulse1, 2 (DPP) beams, such as those used for FLASH radiotherapy3. In these beams, with DPPs from 0.6 Gy to 10 Gy, the response of the ICs is no longer proportional to the dose due to recombination between opposite charge carriers, compromising their usability as reference detectors. In the past years, several numerical and phenomenological models have been developed with the aim of describing the response of parallel-plate ICs under ultra-high DPPs1, 4, 5, 6, 7, 8. These models have shown satisfactory agreement in terms of charge collection efficiency (CCE)5, 9, 10, 11, 4 and time-resolved current4. For parallel-plate ICs, a 1D model assuming cylindrical symmetry around the axis perpendicular to the electrodes provides a close approximation to the actual geometry. However, the picture is different for thimble ICs, where an idealized 1D cylindrical symmetry is likely not to reproduce the actual behavior of the chamber due to its particular combination of cylindrical and spherical geometries. This is particularly dramatic for ICs where the volume of the spherical tip represents a large fraction of the total volume of the IC.
@@
In the literature, the work devoted to volume recombination in cylindrical and spherical ICs is limited. Recently, for example, an ideal 1D cylindrical simulation code was applied to describe the behavior of a PinPoint T31010 IC for DPPs below 0.4 Gy per pulse12. Fenwick et al.13 developed an analytical model including the free electron fraction for cylindrical IC. In ultra-high DPP, only the response of Nano Razor14 and Extradin A26 chambers10 has been experimentally characterized.
@@
II. Materials and methods
II.A. Investigated chambers
This study examines two different models of thimble ICs with similar sensitive volume but different design: the PTW PinPoint 3D T31022 and the PTW PinPoint T31023, hereafter referred to as PP3D and PP, respectively. For each chamber model, two units were assessed—one with internal biasing and the other with external biasing. The key specifications of both IC models are summarized in Table 1. Calibration coefficients in terms of absorbed dose to water were determined at PTB’s 60Co source, using a bias voltage of 400 V. The relative uncertainty of these measurements was 0.4 % (coverage factor ). Although both chambers share similar materials and construction, they differ in their geometric dimensions.
@@ II.A. Investigated chambers
For the 1D and simplified 2D models, the physical volume of the IC was employed to calculate the released charge. In contrast, simulations using the full 2D geometry –which includes the guard ring electrode, accounting for the IC’s dead volume15, 16– utilized the actual sensitive volume of the IC, determined using the simulated electric field lines ending in the central electrode.
II.B. Experimental setup
@@
The experimental characterization of the ICs was conducted at the metrological electron accelerator facility of the German national metrology institute, the Physikalisch-Technische Bundesanstalt (PTB) using the ultra-high pulse dose rate reference electron beam17, 18.
@@
The ICs were irradiated using an electron beam with an energy of 20 MeV. A wide range of DPPs was achieved using different combination of source-to-surface distance (SSD) and aluminum scattering foils as well as a variable diaphragm at the beginning of the beamline to reduce the charge per pulse. Overall, 5 different setups were used: 70 cm SSD (without and with 1 mm Al) and 90 cm SSD (without and with, 1 mm, 2 mm and 6 mm of Al). Two pulse durations were considered in this investigation, 1.0 s and 1.9 s. The chambers were positioned in a PMMA water phantom at a fixed water equivalent depth of 45.2 mm, close to the reference depth. The reference dosimetry was performed using a flashDiamond19, 20 (PTW, Freiburg, Germany), calibrated against alanine pellets21, 22, 11. Pulse-to-pulse deviations produced by variations in the beam charge per pulse were corrected using the signal of an in-flange beam current transformer (ICT)23 (Bergoz, Saint Genis Pouilly, France). The ICT was calibrated with the flashDiamond for each setup, pulse duration and diaphragm aperture to determine the reference DPP. The bias voltage was supplied to the ICs using an in-house high voltage source, varying from 200 V up to 500 V in steps of 100 V for both positive and negative polarity. The charge per pulse from the ICs was measured using a Keythley 616 electrometer with a 33 nF capacitor placed between the core wire and the inner shield of the triaxial cable. This capacitor prevents a non-linear response of the electrometer due to a large signal coming from the detector, as reported by several authors2, 19, 24.
@@
To measure the time-resolved current from the ICs, the collecting electrode was connected to a HVA-200M-40-B voltage amplifier (FEMTO, Berlin, Germany), while the bias voltage was applied to the external electrode of the IC. Simultaneously, the beam pulse structure was monitored using the ICT. Both the IC and ICT currents were recorded simultaneously using the two input channels of a fast 14 bit digitizer (Spectrum M3i.4142), following a procedure similar to that described by Paz-Martín et al.4. Time-resolved measurements were performed using bias voltages ranging from 75 V to 500 V.
@@
II.C. Charge collection efficiency and the polarity effect correction factor determination
@@
The averaged CCE, denoted as in the equations, was determined experimental as the average between the positive and the negative CCE () at the same DPP. was determined as the ratio between the DPP obtained with the ICs (without applying any correction for volume recombination and polarity) over the reference DPP ():
@@
The beam quality correction factors () were calculated using the EGSnrc Monte Carlo code system 25 (version 4.0, 2023a). The linear accelerator beam line and exit window were modeled using the BEAMnrc26 user code, adopting Bourgouin et al.18 approach. To compute the ICs’ perturbation factors, the egs_chamber 27 user code was used, implementing the detailed geometries provided by the ICs’ manufacturer (PTW). The 60Co calibration beam was modeled as a square 10 cm 10 cm divergent beam with a source-to-surface distance of 100 cm and the same energy fluence spectrum as Mora et al. 28. Both electrons and photons were transported until their kinetic energy fell below 1 keV. For the reference electron beam simulations, a cut-off of 10 keV for both electrons and photons was used, as no significant effect was previously reported 29. All the were obtained with a relative statistical uncertainty below 0.35 %.
@@
II.D. The numerical model
The transport equations that describe the drift of the charge carriers inside the ICs were adapted from Paz-Martín et al.4 to describe geometries with cylindrical and spherical symmetry. The system of differential equations considered are:
@@ II.D. The numerical model
To develop a general code capable of simulating a wide variety of geometries, the finite element method was used instead of the traditional finite difference scheme employed for parallel-plate ICs. The code was implemented using the Python package dolfinx 32 (version 0.9.0). To generate the 2D meshes from the IC geometries, the GMSH program 33 was used. The temporal discretization of the equations was performed using an adaptive second-order backward differentiation scheme 34. The numerical code is available at https://github.com/JPazMartin/ICSimFEM.
@@
| Electron velocity | Calculated using Magboltz36 |
| Electron attachment rate | Calculated using Magboltz36 |
| Ion mobilities | From Zhang et al.37 and Boissonnat et al.38 |
| Ion-ion recombination coefficient | Depending on the ion mobilities: for Zhang et al.37 and for Boissonnat et al.38, respectively. |
| Electron-ion recombination coefficient | from Gotz et al.6 (when using Boissonnat et al. mobilities). |
| Longitudinal diffusion coefficient | Calculated using Magboltz36 for electrons and Nernst-Townsend relation for positive and negative ions. |
| Electron multiplication coefficient | Calculated using Magboltz36 |
@@
The time-resolved signal produced by each charge carrier drifting in the IC (denoted as , being positive ions, negative ions, or electrons) was calculated using the Shockley–Ramo theorem 35:
where denotes the actual drift velocity of the charge carriers, and represents the weighting electric field. This field is computed under the following conditions: all free charges are removed from the volume, the collecting electrode set to 1 V and rest of the electrodes grounded. To enable comparison between experimental and simulated time-resolved current signals, the actual beam pulse structure was incorporated into the simulation. The electron and ion transport parameters employed in this study are listed in Table 3.
@@
II.E. The Fenwick et al. analytical model
Recently, Fenwick et al. extended the analytical model developed for calculating the CCE of parallel-plate ICs13, in the absence of electric field perturbation, to describe the CCE of cylindrical ICs40. In their model, the CCE () can be written as:
@@
The parameter has been described previously in Boag’s recombination models41. For a parallel-plate IC with a distance between electrodes , the geometrical parameter is:
@@
III. Results
III.A. Beam quality correction factors
The beam quality correction factors () calculated using EGSnrc for the PP3D and PP ICs are shown in Table 4. The observed variation in across different setups is consistent with changes in the water-to-air stopping power ratio, which result from the degradation of the effective electron energy with aluminum scattering foils of different thickness 39.
@@ III.A. Beam quality correction factors
III.B. Comparison between numerical and analytical model
To initially validate the developed numerical model, the simulated CCE of an idealized cylindrical 1D IC with similar geometrical properties to the PP3D IC was compared to the results of Fenwick et al.40 analytical model (Equation 7). For this comparison, the same transport parameters and simplifications considered by Fenwick et al.40 were incorporated in the numerical model, namely: electric field perturbation was disabled, charge was released instantaneously within the active volume (zero pulse duration), and electron drift was assumed to occur instantaneously after the release of the charge. Figure 2 shows the CCE obtained from the numerical and analytical models as a function of the DPP (from 1 mGy to 10 Gy per pulse) using a bias voltage of 300 V. Discrepancies between the simplified numerical simulation and the analytical values were always below 0.01 % for the whole range of DPPs. For completeness, the results of the complete numerical simulation enabling all of the aforementioned effects are also shown.

### v2

@@ Abstract
Background: Commercially available ionization chambers (ICs) exposed to ultra-high dose per pulse (UHDP) exhibit deviations from a linear dose response due to volume recombination. Simulation models have been developed to describe the charge collection efficiency (CCE) but focused on parallel-plate ICs. This study aims to measure and simulate the CCE and polarity effect of thimble ICs in UHDP.
@@
Methods: The response of two PinPoint3D T31022 (PP3D) and two PinPoint T31023 (PP) ICs was investigated experimentally at the national metrology institute of Germany (PTB). The ICs were irradiated using the UHDP reference electron beam with dose per pulse up to 9.3 Gy for different voltages. A novel finite element code capable of simulating 1D and 2D geometries was developed.
@@
Results: Thimble ICs exhibit a pronounced polarization effect when irradiated with UHDP. When the sign of the collected charge is negative, the CCE is larger than when it is positive. The major contribution to the polarity effect can be attributed to the polarity-dependent charge transport and recombination. Experimental and simulated CCE (polarity effect correction factor) agrees within 1.4 % (7.0 %) and 1.6 % (3.2 %) for the PP3D and PP ICs, respectively. The CCE of parallel-plate and thimble ICs is related through a geometrical rule.
Conclusions: The PP IC shows greater CCE due to its smaller external radius. The numerical model is able to satisfactory reproduce the actual CCE and polarity effect for these two chambers. At UHDP, thimble ICs should be used with caution due to their large polarity effect.
keywords:
Thimble ionization chambers , Charge collection efficiency , Volume recombination , Ultra-high dose rate[USC]organization=Universidade de Santiago de Compostela, addressline=Praza do obradoiro, 0, city=Santiago de Compostela, postcode=15705, state=Galicia, country=Spain
[PTB]organization=Physikalisch-Technische Bundesanstalt, addressline=Bundesallee 100, city=Braunschweig, postcode=38116, country=(Germany)
[IDIS]organization=Group of Medical Physics and Biomathematics, Instituto de Investigacion Sanitaria de Santiago (IDIS), addressline=Travesía da Choupana, city=Santiago de Compostela, state=A Coruña, postcode=15706, country=(Spain)
[CHUS]organization=Department of Medical Physics, Complexo Hospitalario Universitario de Santiago de Compostela, addressline=Travesía da Choupana, city=Santiago de Compostela, state=A Coruña, postcode=15706, country=(Spain)
[RPL]organization=Laboratorio de Radiofísica, Universidade de Santiago, addressline=Estrada de San Lourenzo, city=Santiago de Compostela, postcode=15782, state=A Coruña, country=(Spain)
1 Introduction
@@
Commercially available ionization chambers (IC) suffer from volume recombination when exposed to ultra-high dose per pulse [1, 2] (DPP) beams, such as those used for FLASH radiotherapy [3]. In these beams, with DPPs from 0.6 Gy to 10 Gy, the response of the ICs is no longer proportional to the dose due to recombination between opposite charge carriers, compromising their suitability as reference detectors. In the past years, several numerical and phenomenological models have been developed with the aim of describing the response of parallel-plate ICs under ultra-high DPPs [1, 4, 5, 6, 7, 8]. These models have shown satisfactory agreement in terms of charge collection efficiency (CCE) [5, 9, 10, 11, 4] and time-resolved current [4]. For parallel-plate ICs, a 1D model assuming cylindrical symmetry around the axis perpendicular to the electrodes provides a close approximation to the actual geometry. However, the picture is different for thimble ICs, where an idealized 1D cylindrical symmetry is likely not to reproduce the actual behavior of the chamber due to its particular combination of cylindrical and spherical geometries. This is particularly dramatic for ICs where the volume of the spherical tip represents a large fraction of the total volume of the IC.
In the literature, the work devoted to volume recombination in cylindrical and spherical
@@
ICs is limited. Recently, for example, an ideal 1D cylindrical simulation code was applied to describe the behavior of a PinPoint T31010 IC for DPPs below 0.4 Gy per pulse [12]. Fenwick et al. [13] developed an analytical model including the free electron fraction for cylindrical IC. In ultra-high DPP, only the response of Nano Razor [14] and Extradin A26 chambers [10] has been experimentally characterized.
@@
2 Materials and methods
2.1 Investigated chambers
This study examines two different models of thimble ICs with similar sensitive volume but different design: the PTW PinPoint 3D T31022 and the PTW PinPoint T31023, hereafter referred to as PP3D and PP, respectively. For each chamber model, two units were assessed—one with internal biasing and the other with external biasing. The key specifications of both IC models are summarized in Table 1.
Calibration coefficients in terms of absorbed dose to water were determined at PTB’s 60Co source, using a bias voltage of 400 V. The relative uncertainty of these measurements was 0.4 % (coverage factor ). Although both chambers share similar materials and construction, they differ in their geometric dimensions.
@@ 2.1 Investigated chambers
For the 1D and simplified 2D models, the physical volume of the IC was employed to calculate the released charge. In contrast, simulations using the full 2D geometry –which includes the guard ring electrode, accounting for the IC’s dead volume [15, 16]– utilized the actual sensitive volume of the IC, determined using the simulated electric field lines ending in the central electrode.
2.2 Experimental setup
@@
The experimental characterization of the ICs was conducted at the metrological electron accelerator facility of the German national metrology institute, the Physikalisch-Technische Bundesanstalt (PTB) using the ultra-high pulse dose rate reference electron beam [17, 18].
@@
The ICs were irradiated using an electron beam with an energy of 20 MeV. A wide range of DPPs was achieved using different combination of source-to-surface distance (SSD) and aluminum scattering foils as well as a variable diaphragm at the beginning of the beamline to reduce the charge per pulse. Overall, 5 different setups were used: 70 cm SSD (without and with 1 mm Al) and 90 cm SSD (without and with, 1 mm, 2 mm and 6 mm of Al). Two pulse durations were considered in this investigation, 1.0 s and 1.9 s. The chambers were positioned in a PMMA water phantom at a fixed water equivalent depth of 45.2 mm, close to the reference depth of the different setups. The reference dosimetry was performed using a flashDiamond [19, 20] (PTW, Freiburg, Germany), calibrated by means of PTB’s alanine dosimetry system [21, 22, 11]. Pulse-to-pulse deviations produced by variations in the beam charge per pulse were corrected using the signal of an in-flange beam current transformer (ICT) [23] (Bergoz, Saint Genis Pouilly, France). The ICT was calibrated with the flashDiamond for each setup, pulse duration and diaphragm aperture to determine the reference DPP. The bias voltage was supplied to the ICs using an in-house high voltage source, varying from 200 V up to 500 V in steps of 100 V for both positive and negative polarity. The charge per pulse from the ICs was measured using a Keythley 616 electrometer with a 33 nF capacitor placed between the core wire and the inner shield of the triaxial cable. This capacitor prevents a non-linear response of the electrometer due to a large signal coming from the detector, as reported by several authors [2, 19, 24].
@@
To measure the time-resolved current from the ICs, the collecting electrode was connected to a HVA-200M-40-B voltage amplifier (FEMTO, Berlin, Germany), while the bias voltage was applied to the external electrode of the IC. Simultaneously, the beam pulse structure was monitored using the ICT. Both the IC and ICT currents were recorded simultaneously using the two input channels of a fast 14 bit digitizer (Spectrum M3i.4142), following a procedure similar to that described by Paz-Martín et al. [4]. Time-resolved measurements were performed using bias voltages ranging from 75 V to 500 V.
@@
2.3 Charge collection efficiency and the polarity effect correction factor determination
@@
The averaged CCE, denoted as in the equations, was determined experimentally as the average between the positive and the negative CCE () at the same DPP. was determined as the ratio between the DPP obtained with the ICs (without applying any correction for volume recombination and polarity) over the reference DPP ():
@@
The beam quality correction factors () were calculated using the EGSnrc Monte Carlo code system [25] (version 4.0, 2023a). The linear accelerator beam line and exit window were modeled using the BEAMnrc [26] user code, adopting Bourgouin et al. [18] approach. To compute the ICs’ perturbation factors, the egs_chamber [27] user code was used, implementing the detailed geometries provided by the ICs’ manufacturer (PTW). The 60Co calibration beam was modeled as a square 10 cm 10 cm divergent beam with a source-to-surface distance of 100 cm and the same energy fluence spectrum as Mora et al. [28]. Both electrons and photons were transported until their kinetic energy fell below 1 keV. For the reference electron beam simulations, a cut-off of 10 keV for both electrons and photons was used, as no significant effect was previously reported [29]. All the simulated were obtained with a relative statistical uncertainty below 0.35 %. Additionally the factors were experimentally determined for two different detectors of each type by cross-calibrating them with a PTW 34001 Roos chamber in an electron beam with a nominal Energy of 20 MeV (R50 = 77.4 mm) produced by a clinical LINAC (Elekta Synergy).
@@
2.4 The numerical model
The transport equations that describe the drift of the charge carriers inside the ICs were adapted from Paz-Martín et al. [4] to describe geometries with cylindrical and spherical symmetry. The system of differential equations considered are:
@@ 2.4 The numerical model
To develop a general code capable of simulating a wide variety of geometries, the finite element method was used instead of the traditional finite difference scheme employed for parallel-plate ICs. The code was implemented using the Python package dolfinx [32] (version 0.9.0). To generate the 2D meshes from the IC geometries, the GMSH program [33] was used. The temporal discretization of the equations was performed using an adaptive second-order backward differentiation scheme [34].
@@
| Electron velocity | Calculated using Magboltz [36] |
| Electron attachment rate | Calculated using Magboltz [36] |
| Ion mobilities | From Zhang et al. [37] and Boissonnat et al. [38] |
| Ion-ion recombination coefficient | Depending on the ion mobilities: for Zhang et al. [37] and for Boissonnat et al. [38], respectively. |
| Electron-ion recombination coefficient | from Gotz et al. [6] (when using Boissonnat et al. mobilities). |
| Longitudinal diffusion coefficient | Calculated using Magboltz [36] for electrons and Nernst-Townsend relation for positive and negative ions. |
| Electron multiplication coefficient | Calculated using Magboltz [36] |
@@
The time-resolved signal produced by each charge carrier drifting in the IC (denoted as , being positive ions, negative ions, or electrons) was calculated using the Shockley–Ramo theorem [35]:
where denotes the actual drift velocity of the charge carriers, and represents the weighting electric field. This field is computed under the following conditions: all free charges are removed from the volume, the collecting electrode set to 1 V and rest of the electrodes grounded.
To enable comparison between experimental and simulated time-resolved current signals, the actual beam pulse structure was incorporated into the simulation. The electron and ion transport parameters employed in this study are listed in Table 3.
@@
2.5 The Fenwick et al. analytical model
Recently, Fenwick et al. extended the analytical model developed for calculating the CCE of parallel-plate ICs [13], in the absence of electric field perturbation, to describe the CCE of cylindrical ICs [40]. In their model, the CCE () can be written as:
@@
The parameter has been described previously in Boag’s recombination models [41]. For a parallel-plate IC with a distance between electrodes , the geometrical parameter is:
@@
3 Results
3.1 Beam quality correction factors
The experimental and calculated beam quality correction factors () for the PP3D and PP ICs are shown in Table 4. The observed variation in across different setups is consistent with changes in the water-to-air stopping power ratio, which result from the electron energy degradation and straggling with aluminum scattering foils of different thickness [39]. The 20 MeV Elekta Synergy beam is comparable in terms of beam quality to the 90 cm + 2 mm Al setup, with an value of 77.1 mm. The experimental determined , averaged between the two ICs samples, show a good agreement with the simulated values.
@@ 3.1 Beam quality correction factors
| 20 MeV Elekta Synergy | 0.9145 | 0.9150 |
3.2 Comparison between numerical and analytical model
To initially validate the developed numerical model, the simulated CCE of an idealized cylindrical 1D IC with similar geometrical properties to the PP3D IC was compared to the results of Fenwick et al. [40] analytical model (Equation 7). For this comparison, the same transport parameters and simplifications considered by Fenwick et al. [40] were incorporated in the numerical model, namely: electric field perturbation was disabled, charge was released instantaneously within the active volume (zero pulse duration), and electron drift was assumed to occur instantaneously after the release of the charge. Figure 2 shows the CCE obtained from the numerical and analytical models as a function of the DPP (from 1 mGy to 10 Gy per pulse) using a bias voltage of 300 V. Discrepancies between the simplified numerical simulation and the analytical values were always below 0.01 % for the whole range of DPPs. For completeness, the results of the complete numerical simulation enabling all of the aforementioned effects are also shown.